The measure of one angle of a right triangle is 3 degrees less than twice the measure of the smallest angle. Find the measure of each angle.
Answer by JBarnum(2146)

90+90=180
THERE IS 180 DEGREES IN EVERY TRIANGLE
SO THE OTHER 2 ANGLES IN A RIGHT TRIAGLE MUST ADD UP TO 90
S=SMALLEST ANGLE
2S-3+S=90
3S-3=90
3S=93
S=31
O=90-31
O=59
SO THE ANGLES ARE 31, 59, 90